# A Plug-and-Play Learning-based IMU Bias Factor for Robust Visual-Inertial Odometry

### v1

@@ III-D IMU-Bias Label Solving Method
$\hat{\gamma}_{b_{k+1}}^{b_{k}}{}^{-1}\textperiodcentered\gamma_{b_{k+1}}^{b_{k}}$ $\approx\left[\frac{1}{2}J_{b_{\omega}}^{\gamma}\delta b_{\omega}\right].$
where the specific form of $J$ refers to [4]. To directly obtain the mean values of $\mathbf{b}_{a_{t}}$ and $\mathbf{b}_{w_{t}}$, similarity constraints can be further applied to iteratively derive the final IMU bias, which can be used as labels for network training.
@@ IV EXPERIMENTS
We conduct experiments on both public datasets and real-world scenarios to evaluate the proposed framework. In the first experiment, we select the combination of monocular vision and IMU from VINS-Fusion [5] as the baseline and conduct comparison studies to validate the localization precision and robustness on two public datasets. Furthermore, we test our framework on a self-collected dataset to assess its performance on low-cost IMU. Additionally, we use two public datasets with IMU bias labels to verify the accuracy of our iterative method.
IV-A Experimental Setup
Figure 5: (a) The image illustrates the specific environment and setup during the data collection process; (b) The image compares the differences between the trajectory generated by the baseline method and our proposed framework with the ground truth trajectory on the Seq02 sequence; (c) The image illustrates the spatial relationship between the IMU and the camera within our data acquisition system.
IV-A1 Datasets
@@
We select the EuRoC [33], TumVi [34], AR_Table [35] and in-house datasets for the evaluation. EuRoC: This dataset is widely used and consists of 20Hz drone images and 200Hz inertial data, with accurate the ground truth pose obtained through a motion capture system. Fortunately, each sub-dataset provides the ground truth of IMU bias, offering a valuable reference for verifying the effectiveness of our bias-solving method.
TumVi: This dataset was collected using handheld devices in various scenarios and is one of the commonly used datasets for evaluating VIO. It provides raw data sequences of two different image resolutions along with corresponding calibration sequences. It is worth noting that the IMU data in the calibration sequences have been corrected for bias. In our experiments, we compensate for this bias based on the parameters provided on the official website.
AR_Table: This dataset focuses on indoor tables as the collection target and aims to verify SLAM algorithms using planar features. Similar to the EuRoC dataset, it also provides accurate the ground truth pose and the ground truth bias through a motion capture system. Due to the dataset’s rich visual features, the IMU plays a limited role in constraints, so it is used solely to evaluate the accuracy of IMU bias computation.
In-house Dataset: To further validate the effectiveness of the framework, we select a low-cost IMU device, WHEELTEC N100, with a sampling frequency of 200Hz, and paired it with the Realsense D456 to form a simple visual-inertial data collection unit. The spatial position relationship of the hardware module is shown in Fig.5(c). We collected 12 data sequences in the environment shown in Fig.5(a), with a cumulative duration of about 20 minutes.
IV-A2 Evaluation Metrics
Firstly, to evaluate localization precision, we use Absolute Trajectory Error (ATE, m) and Relative Translational Error (RTE, m) as metrics, taking each sub-dataset as the smallest unit of evaluation, and select Root Mean Squared Error (RMSE) as the statistical basis. Secondly, to evaluate localization robustness, we describe it intuitively through the divergence and convergence of the trajectory. Finally, to evaluate the accuracy of the iterative method for calculating IMU bias, we compare the calculated results with the average values of each dimension in the sub-dataset individually.
IV-A3 Implementation Details
@@
We implement the entire framework on a Lenovo Y9000P computer equipped with an RTX 4090 GPU, running Ubuntu 20.04, and using OpenCV 3.4.5.
@@
The proposed IPNet is designed and implemented using the PyTorch framework, with MAE as the loss function. The RMSprop optimizer is employed with an initial learning rate of 1e-6, which is reduced by a factor of 10 after every 10 iterations. Other hyperparameter settings are provided in the source code. During the training of IPNet, we conduct a statistical analysis of the sample distribution in each sub-dataset and observe an imbalance in the data distribution. To address this, we adopt a strategy different from previous studies [22, 6, 24]. For the Euroc dataset, MH_01, MH_03, MH_04, V1_03, and V2_03 are used as the training set, MH_02 and V1_01 as the validation set, and the remaining data as the test set. For the TumVI dataset, room1, room2, and room4 are selected as the training set, room3 as the validation set, and the remaining two datasets as the test set. For the self-collected dataset, six sequences are chosen for the training set, two for the validation set, and the remaining four sequences for the test set.
We implement IMU bias estimation in PyTorch using a two-stage approach to obtain bias labels for network training. In the first stage, we calculate the bias of the gyroscope using the Adam optimizer with a learning rate of 0.001, 15000 iterations, and a 10-fold reduction in the learning rate every 5000 iterations. In the second stage, the gyroscope bias is fixed, and the bias of the accelerometer is calculated. The hyperparameters remain the same except for the learning rate, which is set to 0.01.
IV-B Performance
Tab. I shows a comparison of the performance between the baseline method and our framework on public datasets. The results demonstrate that our method achieves significant improvements, with the ATE-RMSE improving by an average of 46%, and the RTE-RMSE improving by an average of 8%. Additionally, the baseline method fails to track the trajectory of sequence V2_02 and room5, leading to divergence. This is primarily due to large visual tracking errors in these sequences, combined with a low image processing frame rate of only 10Hz. The original inertial constraints are insufficient to compensate for the system, thus causing trajectory divergence. Taking the room5 sequence as an example, frequent shaking leads to drastic changes in the viewpoint, coupled with the limitation on the minimum distance for visual matching, which results in a limited number of feature points that can be correctly used at certain moments, ultimately causing the estimate to diverge. In contrast, our method introduces IMU bias prior, providing relatively accurate IMU constraints to the system, significantly improving the localization precision and robustness, even in situations with large visual errors.
Figure 6: IMU Bias Calculation Accuracy Experiment. The horizontal axis represents the ground truth bias, and the vertical axis represents the fake label of bias. Ideally, in the ideal case, they should be equal and distributed along a straight line with a slope of 1 and passing through the origin. The figure displays the bias distribution and fitting of different datasets. The gray area represents the confidence interval, used to characterize the 95% confidence interval of the fitting curve, while the yellow area represents the prediction interval, used to characterize the estimated range of future observation values under the 95% confidence level.
TABLE I: ATE [M] & RTE [M]: Performance on public datasets
@@ IV-B Performance
ATE RTE ATE RTE
Euroc MH_05 0.340 0.0074 0.249(27%) 0.0065(12%)
V1_02 0.270 0.0101 0.106(61%) 0.0074(26%)
V2_01 0.133 0.0138 0.048(64%) 0.0066(52%)
V2_02 - - 0.098 0.0055
TumVI room5 - - 0.436 0.1044
room6 0.198 0.0352 0.102(48%) 0.0403(-15%)
Average 0.235 0.0166 0.126(46%) 0.0152(8%)
@@
Fig.4 shows the variation of the bias corresponding to each pose during the VIO tracking process for the baseline and our methods in the V1_02 sequence. By introducing the prior, it is clearly observed that the bias are effectively constrained within a reasonable range. This approach not only improves the accuracy of bias estimation but also makes the optimization process more consistent with real physical constraints, thereby enhancing the overall stability and reliability of the optimization.
To facilitate a comparison with previous recursive prediction methods, we adopt the same training strategy for our network and present the final localization results in Tab. II. The results show that our method outperforms the previous approaches.
TABLE II: ATE [M]: Performance on Euroc dataset
@@
MH_02 0.150 0.130 0.087
MH_04 0.138 0.250 0.188
V1_01 0.066 0.080 0.053
V1_03 0.147 0.170 0.147
V2_02 0.104 0.100 0.081
Tab. III shows the comparison results between our method and the baseline method on the self-collected dataset. Among them, Seq04 is purposely selected as a scene with high speed and drastic changes in the field of view. In this sequence, the original baseline on which the trajectories rapidly diverge, while our method can effectively maintain the tracking state and improve the robustness of the system. For the remaining sub-datasets, our method shows significant improvements in localization precision compared to the baseline, with the ATE-RMSE improving by an average of 46%, of which the trajectory comparison effect graph of the Seq02 sequence is shown in Fig.5(b). Due to the limitations in the number and layout of Vicon systems in the experimental environment, we deliberately reduce the rotation of the rigid body during data collection to prevent frequent loss of ground truth data. As a result, no significant improvement is observed in the RTE-RMSE metric during testing. Overall, the experimental results are consistent with our findings on public datasets.
TABLE III: ATE [M] & RTE [M]: Performance on Our Dataset
@@
IV-C IMU Bias Calculations VS. Labels
We validate the accuracy of our iterative method on the Euroc and AR_Table datasets and compare the calculated results with the ground truth average. The comparison results are shown in Fig.6. The slope and intercept of the curves in the figure clearly reflect the accuracy of the calculation, which can be directly used as labels for network training.
V CONCLUSIONS
In this paper, we propose IPNet, which directly estimates IMU bias from raw data in real time and integrates bias estimates as a plug-and-play prior factor into the VIO framework. The framework effectively reduces IMU bias estimation errors arising from unreliable visual features, thereby significantly improving localization accuracy and enhancing system robustness. The framework was implemented and evaluated across multiple datasets, with experimental results demonstrating that precise bias estimation leads to substantial improvements in overall system performance. In the near future, we aim to further optimize network training by incorporating noise labels for fine-tuning, ensuring that the estimations align more accurately and comprehensively with the IMU noise model.

### v2

@@ III-D IMU bias estimation for training label
$\hat{\gamma}_{b_{k+1}}^{b_{k}}{}^{-1}\cdot\gamma_{b_{k+1}}^{b_{k}}$ $\approx\left[\frac{1}{2}J_{b_{\omega}}^{\gamma}\delta b_{\omega}\right].$
where the specific form of $J$ refers to [16]. Assuming the initial values of $b_{a_{init}}$ and $b_{w_{init}}$ are 0, a linear expansion at this point can be used to calculate the bias at each moment as $\mathbf{b}{a_{t}}=\delta b_{a}+b_{a_{init}}$ and $\mathbf{b}{w_{t}}=\delta b_{w}+b_{w_{init}}$. To directly obtain the mean values of $\mathbf{b}{a_{t}}$ and $\mathbf{b}{w_{t}}$, a similarity constraint can be further applied to the entire sequence, iteratively solving for the final IMU bias. These biases can then be used as the training labels for the network.
@@ IV Experiments
Two public datasets with IMU bias labels were used to verify the accuracy of our iterative method and its resulting IMU bias labels. The resulting label data is released publicly. Then, we evaluated the proposed framework on public datasets and In-house dataset. Localization precision and robustness were validated against the VINS-Fusion[5] Mono-IMU baseline across two public datasets. Finally, performance assessment was also conducted on In-house dataset, followed by generalization testing through cross-domain transfer of indoor-trained models to outdoor environments.
IV-A Accuracy evaluation of IMU bias calculations in training set and training details
We validate the accuracy of our iterative method on the EuRoC and AR_Table datasets by comparing the calculated results with the ground truth average. The comparison results are shown in Fig.4. The slope and intercept of the curves in the figure clearly reflect the calculation accuracy, which can be directly used as labels for network training.
IV-B Experimental setup
Figure 5: The image shows schematic diagrams of different scenes in challenging environments, which correspond to the following test sequences: (a): the Seq04 sequence of the In-house dataset, (b): the Seq17 sequence of the In-house dataset, (c): the V2_02 sequence of the EuRoC dataset, and (d): the room_5 sequence of the TumVi dataset. Due to poor visual features in these scenes, the trajectories diverge, affecting the robustness of the system.
IV-B1 Datasets
We select the EuRoC [35], TumVi [36], AR_Table [37] and In-house dataset for the evaluation.
@@ IV-B1 Datasets
EuRoC: This dataset is widely used and consists of 20Hz drone images and 200Hz inertial data, with accurate the ground truth pose obtained through a motion capture system. Fortunately, each sub-dataset provides the ground truth of IMU bias, offering a valuable reference for verifying the effectiveness of our bias-solving method.
TumVi: This dataset was collected via handheld platforms across diverse scenarios, exhibiting substantially different motion patterns from the EuRoC’s smooth aerial maneuvers, thus serving as an established benchmark for evaluating VIO system robustness.
AR_Table: This dataset is primarily designed for 3D reconstruction, providing motion-capture ground truth poses and biases similar to EuRoC. Due to its visually rich scenes diminishing the reliance on IMU constraints, it is exclusively used for evaluating IMU bias estimation accuracy.
In-house Dataset: We built a modular visual-inertial sensor unit using the low-cost WHEELTEC N100 IMU (200Hz) and Realsense D456 camera, as shown in Fig. 6. Initially, the unit was integrated into a handheld device, and 12 motion sequences were collected indoors (total: 20 minutes). Since only one side of the Vicon system was functional, we focused on planar motion to ensure reliable pose ground truth. The unit was then transferred to an autonomous vehicle platform, where 5 additional sequences were collected outdoors (total: 8 minutes), with pose ground truth obtained using the GPS mode in VINS-Fusion [5]. The GPS module we use is the UBLOX ZED-F9P, with an RTK positioning accuracy of 0.01m + 1ppm CEP.
Fig. 5 shows several challenging scenes from the dataset. Due to issues such as motion blur, sparse textures, and low lighting, the visual features are poor for a common VIO system. As a result, the baseline system fails to localize in these scenes.
Figure 6: The image respectively illustrates the environmental setup, equipment platform, and partial trajectory comparison results obtained during the data collection process. The top trajectory corresponds to the Seq02 sequence of the in-house dataset, while the bottom trajectory corresponds to the Seq16 sequence of the same dataset.
IV-B2 Evaluation metrics
Localization precision is quantified using absolute trajectory error (ATE-RMSE/M) and relative pose error (RPE-RMSE/RAD) per sub-dataset. System robustness is characterized visually through trajectory divergence/convergence phenomena. IMU bias accuracy is validated by comparing iterative solutions against per-dimension bias means.
IV-B3 Implementation details
@@
Experiments were conducted on a Lenovo Y9000P platform equipped with an RTX 4090 GPU, running Ubuntu 20.04 with OpenCV 3.4.5.
@@
We implemented the proposed IPNet using the PyTorch framework, with MAE as the loss function and RMSprop as the optimizer. The initial learning rate is set to 1e-6 and decays by a factor of 10 every 10 epochs. Statistical analysis of the sample distribution revealed significant imbalance across sub-datasets. To address this, we adopted a training strategy different from previous studies [26, 7, 28].The dataset splits are as follows: for the EuRoC dataset, MH_01, MH_03, MH_04, V1_03, and V2_03 are used for training, MH_02 and V1_01 for validation, and the remaining sequences for testing; for TumVi, room1, room2, and room4 are used for training, room3 for validation, and the remaining two sequences for testing; for In-house Dataset, six indoor sequences are used for training, two indoor sequences for validation, and the remaining four indoor sequences together with all outdoor sequences form the test set. Finally, to ensure a fair comparison with previous recursive methods, we conducted additional experiments by retraining those models under the same data splitting strategy for evaluation.
TABLE I: ATE [M] & RPE [RAD]: Performance on public datasets
@@
ATE RPE ATE RPE
EuRoC MH_05 0.340 0.02115 0.249(27%) 0.00655(69%)
V1_02 0.270 0.03022 0.106(61%) 0.00754(75%)
V2_01 0.133 0.04317 0.048(64%) 0.00672(84%)
V2_02 - - 0.098 0.00565
TumVi room5 - - 0.436 0.10512
room6 0.198 0.03645 0.102(48%) 0.04145(-14%)
Average 0.235 0.0327 0.126(46%) 0.0157(48%)
@@
- indicates trajectory divergence of the baseline.
For the calculation of IMU bias labels, we first optimize the gyroscope bias using the Adam optimizer with an initial learning rate of 0.001 for 15,000 iterations, reducing the learning rate by a factor of 10 every 5,000 iterations. The resulting gyroscope bias is then fixed, and the accelerometer bias is optimized with a learning rate of 0.01, while keeping all other hyperparameters unchanged.
IV-C Performance
Figure 7: The image shows the trajectory and pose comparison between our method and the baseline for the V2_01 sequence of the EuRoC dataset, with the red box highlighting the comparison during the initial phase.
Figure 8: The comparison of bias values curves corresponding to each pose in the V1_02 sequence of the EuRoC dataset. In the baseline, the bias values corresponding to the first 200 poses deviate significantly from the ground truth, whereas with our method, the bias values are quickly adjusted to fluctuate close to the ground truth. Pink: IPNet output; Blue: the bias predicted by baseline; Green: the bias predicted by our framework; Red: our label.
As shown in Table I, the proposed framework significantly enhances localization precision on public datasets, with ATE-RMSE and RPE-RMSE improved by an average of 46% and 48%, respectively. The baseline method experiences trajectory divergence in sequences V2_02 and room5, mainly due to large visual tracking errors and a low image processing rate of only 10Hz, which leads to insufficient inertial constraints for error compensation. In the case of room5, frequent shaking causes drastic changes in viewpoint, and the minimum distance threshold in visual matching limits the number of usable feature points at certain moments, eventually resulting in divergence. In contrast, our method introduces an IMU bias prior, providing more accurate bias constraints to the system, thereby significantly enhancing localization accuracy and robustness even under severe visual degradation. It is worth noting that in the room6 sequence, the RPE-RMSE slightly decreased, possibly due to the limited training data and the simplicity of the motion pattern, which prevented the network from learning more complex motion patterns.
During the VIO initialization phase, the system is often affected by poor visual features and severe vibrations, which cause the estimated IMU bias to deviate significantly from the actual physical value, thus severely affecting the accuracy of the initial pose. This deviation can lead to a decline in the overall performance. As shown in Fig. 7, during the initialization phase, our approach applies constraints during the IMU bias estimation process, effectively limiting its deviation and ensuring the accuracy of both IMU bias and pose estimation.
Fig. 8 shows the pose bias variations during VIO tracking for our method versus the baseline on sequence V1_02. By incorporating prior, biases are effectively constrained within reasonable bounds, enhancing estimation accuracy and system stability. While the gyroscope’s Z-axis estimates are marginally inferior to the baseline due to shifted optimization priorities following prior introduction, their numerical discrepancy remains minimal.
TABLE II: ATE [M] & RPE [RAD]: Performance on EuRoC dataset
@@ IV-C Performance
ATE RPE ATE RPE ATE RPE
MH_02 0.150 0.0025 0.130 - 0.087 0.0026
MH_04 0.138 0.0037 0.250 - 0.188 0.0072
V1_01 0.066 0.0162 0.080 - 0.053 0.0056
V1_03 0.147 0.0067 0.170 - 0.147 0.0110
V2_02 0.104 0.0068 0.100 - 0.081 0.0044
- RPE not reported in the original paper.
To facilitate direct comparison with recursive methods, we retrained and evaluated our model using identical training protocols (Table II). While our approach demonstrates measurable advantages, fundamental architectural disparities preclude strictly equivalent comparison. Nevertheless, the experimental results show that the non-recursive method for learning a stable bias prior is highly competitive with the recursive time-varying bias estimation framework.
TABLE III: ATE [M] & RPE [RAD]: Performance on In-house Dataset
@@
indicates that the sequence was collected in the evening.
✓indicates that the trajectory converges(lacking ground truth).
- indicates trajectory divergence of the baseline.
Tab. III shows the comparison results between our method and the baseline method on the self-collected dataset. In terms of localization accuracy, the proposed method improves the ATE-RMSE by an average of 20% compared to the baseline, with a more significant improvement in indoor scenes than outdoor ones, which aligns with expectations. However, it did not show a significant advantage in the RPE-RMSE metric. This can be attributed to two main factors: first, the limited coverage of the Vicon system in indoor data collection, which focused mainly on planar motion, leading to insufficient learning of roll and pitch modeling; second, the presence of dynamic objects, such as pedestrians and vehicles, in outdoor scenes, which affected pose estimation accuracy. For Seq17, we deliberately selected an evening scene with significant visual errors. Due to the divergence of the baseline trajectory, accurate pose ground truth could not be obtained, and thus, quantitative results cannot be provided. Overall, while improvements in RPE-RMSE are limited, the method demonstrates notable advantages in system robustness, which is consistent with our previous findings on public datasets.
V Discussion
V-A Analysing of Inference Speed
As shown in Fig. 9, We used a subset of the EuRoC dataset to calculate the average inference time of the network, with a total of over 25,000 inference iterations. The results show that the IPNet achieves an inference frequency of 500Hz, which is 2.5 times the IMU output frequency. Combined with the time for feature point tracking and backend optimization solving, the overall system can achieve a running frequency of 30Hz, meeting the real-time operation requirements. The entire real-time performance test was conducted on a system equipped with an RTX 4090 GPU running Ubuntu 20.04.
Figure 9: The time performance analysis in the MH_03 sequence of the EuRoC dataset.
V-B Analysing of Generalization of IPNet
We performed data collection using a handheld device in an indoor environment, then completed model training and evaluation. Subsequently, we applied the trained model to the outdoor autonomous vehicle equipment. As shown in the experimental results in Tab. III, although there was an improvement in positioning accuracy and robustness, and the model demonstrated good generalization ability, the improvement was slightly reduced compared to the indoor environment. The evaluation demonstrated the generalization of the proposed IPNet on different ground platforms with similar motion patterns.
V-C Analysing the requirement of Vicon
In this work, we collect GT pose using Vicon for calculating the bias labels. Then, we come up with a question: can we get the valid labels without Vicon, so scale up the application of the proposed method?
To answer this question, we conducted the following experiments. Fig. 10 shows the difference between the IMU biases directly obtained by VIO and those obtained by Vicon in the EuRoC dataset. In the MH_01 and MH_02 sequences, the differences are relatively small, but in other sequences, the differences become very significant due to the substantial increase in visual errors.
From the experiments, we can observe that good visual features can help the VIO backend estimate the IMU bias accurately. Since our network uses the average value of the IMU bias as the supervision signal, and the average naturally smooths out noise. In the cases with good visual features, our method can obtain the IMU bias iteratively using the pose generated by VIO, without relying on the ground truth pose, and use it as the supervision signal. However, in scenarios with significant visual errors, it is necessary to rely on Vicon to obtain ground truth for iterative solving. This insight paves the way for our future work: building a scalable, self-supervised training pipeline by automatically curating high-confidence segments from large-scale VIO trajectories, effectively eliminating the dependency on external motion capture systems.
Figure 10: The difference between the results obtained by directly solving for the bias using the poses generated by VIO and those obtained using Vicon. (a): The differences in the three axes of the gyroscope, (b): The differences in the three axes of the accelerometer.
VI Conclusions
In this paper, we propose IPNet, a method that directly infers IMU bias prior from raw IMU data using a sliding window approach. The inferred prior is integrated into the VIO framework as plug-and-play factors. This effectively reduces IMU bias estimation errors caused by unreliable visual features, significantly improving positioning accuracy and enhancing the system’s robustness. We implemented and evaluated the module on multiple datasets. Experimental results show that while the bias prior estimates output by the network are not the final precise values, they can significantly improve the overall system performance through collaborative optimization with the VIO backend. Future plans include introducing noise label optimization for training, and based on the strong robustness of the current device-specific strategy, further exploring selecting the good sequence automatically to make the proposed method scalable without Vicon.